OpenStudy (anonymous):

a drawer contains four red socks and five blue socks. Three socks are dawn. Determine the probability that a) exactly 2 red socks are selected b) at least 2 red socks are drawn

OpenStudy (kropot72):

\[P(2\ red)=\frac{4C2 \times 5C1}{9C3}\]

OpenStudy (kropot72):

\[P(3\ red)=\frac{4C3}{9C3}\] The probability that at least 2 red socks are drawn is P(2 or more red) = P(2 red) + P(3 red)

OpenStudy (ybarrap):

P(a)=(3 2 )4∗3∗2/(9∗8∗7) P(b)=P(a)+(3 3 )∗4∗3∗5/(9∗8∗7) (both our solutions should give the same answer, 5/14 and 1/21)